The quality assurance department at a hospital has as one of its responsibilities the verification of medical records. One of the department’s employees samples medical records and checks each one for correctness. Historical data suggests that each medical record has a 0.98 probability of being correct. Suppose the employee begins checking medical records at some point in time. Let X be a random variable denoting whether a single medical record being inspected is correct or not. Determine what type of random variable X is.
Answer #1

TOPIC:Bernoulli distribution.

* Define, Þ= P (of a selected medical record is connect) = 0.98 q =P (of a selected medical recond is inconnect). So, = 1-p (